The capital structure of ABC Pvt. Ltd is as follows: Equity share capital (eachshareofRs.10) = Rs.16, 00,000Debentures with a coupon rate of 10% = Rs. 10, 00,000Reserves and surplus = Rs.15, 00,000Revenue from the business activities for the company is Rs. 2.00 crores. Its variable cost is 10% of the revenue, fixed operating cost is Rs. 60 lakhs and the company pays income tax at a rate of 25%. (10Marks)Calculate financial leverage, operating leverage and combined leverage for the company.

1
Expert's answer
2021-08-30T14:32:55-0400

Solution:

a.). Financial Leverage ="\frac{Total \; Debt}{Shareholders\; Equity}"

Total Debt = 1000,000

Shareholders Equity = Equity shares + Reserves and Surplus = 1,600,000 + 1,500,000 = 3,100,000


Financial Leverage = "\frac{1,000,000}{3,100,000} = 0.32"


Financial Leverage = 0.32

 

Operating Leverage = "\frac{Fixed \; Costs}{Total\; Costs}"


Fixed Costs = 60 lakhs = 0.60 crores

Variable Costs = "10\%\times 2" crores = 0.20 crores

Total Costs = FC + VC = 0.60 + 0.20 = 0.80 crores


Operating Leverage = "\frac{0.60}{0.80} = 0.75"


Operating Leverage = 0.75

 

Combined Leverage = Operating Leverage x Financial Leverage = 0.75 "\times" 0.32 = 0.24

 

Combined Leverage = 0.24